OpenStudy (anonymous):

If f(x) = sin(x) for all x, then the average value of f on the interval [0, π] is (A) ½ (B) 1/π (C) π/2 (D) 2/pie

OpenStudy (anonymous):

pie <- yummeh

OpenStudy (anonymous):

Are you taking calc

OpenStudy (anonymous):

yes Download: www.ieType.com/f.php?FrOyOg

OpenStudy (anonymous):

no, I mean are you in calculus course?

OpenStudy (anonymous):

i dont know hw to cal!

OpenStudy (anonymous):

\[(1/\pi)\int\limits_{0}^{\pi}\sin(x)\]

OpenStudy (anonymous):

what hapnd guys

OpenStudy (anonymous):

dx* my bad sorry

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Well the average value of a function is represented by (1/b-a) times the integral of the function with a as the lower limit and b as the upper limit. Using the fundamental theorm of calculus to solve the integral from a to b of f(x) you would find F(x), the antiderivative of f(x), and then do F(b)-F(a) to solve for the integral. In this case, the antiderivative would be -cos(x) because the derivative of -cos(x) is sin(x). So, -cos(pi) + cos(0) would be solved to get 2, and then multiplied by (1/pi), which would get you the answer of (2/pi), or D. Sorry for saying B I meant D I just did the integral quickly in my head and messed up, still new at this. Answer: D